If each of two opposite sides of a square is increased by 3 centimetres and each of the other two sides is decreased by 2 centimetres, the area is increased by 8 square centimetres. Find the length of a side of the square.
Answer by lwsshak3(11628)

If each of two opposite sides of a square is increased by 3 centimetres and each of the other two sides is decreased by 2 centimetres, the area is increased by 8 square centimetres. Find the length of a side of the square.
***
let x=length of a side of the square.
..
area of a rectangle=length*width
(x+3)(x-2)-x^2=8
x^2+x-6-x^2=8
x-6=8
x=14
length of a side of the square=14 cm